# Milliarcsecond astrometric oscillations in AGN as a precursor of multi-messaging gravitational wave events

### v1

@@ Abstract
Aims. Previously we have identified the quasar J21026015 as potential SMBHB system on the basis of absolute astrometric very long baseline interferometry (VLBI) monitoring. Here we present another case, the source J02041514, exhibiting similar oscillating astrometric pattern. Our aim is to analyse evolution of SMBHB as a generator of GW and provide physical ‘multi-messaging’ link between astrometric manifestation in radio domain and GW emission.
Methods. We analysed the available in the archives VLBI astrometry data that resulted in detection of astrometric oscillations in the source J02041514. We assume these oscillations as manifestations of orbital motion in a binary systems. We estimated parameters of the suspected SMBHB in this source and applied basic theoretical models for projecting its evolution toward coalescence. We also develop a toy model of SMBHB consistent with the discovered astrometric oscillations and give quantitative predictions of GW emission of such the source using the case of J02041514 as an example potentially applicable to other SMBHBs.
Results. We provide observational evidence of astrometric oscillations in the source J02041514. As an ad hoc result, we also provide a re-assessed estimate of the redshift of J21026015, . A toy model of the object containing SMBHB with parameters consistent with the observed astrometric oscillations of the source J02041514 as an example enabled us to consider GW emission as the cause of the system’s orbital evolution.
Conclusions. We conclude that astrometric VLBI monitoring has an appreciable potential for future detections of SMBHBs which could become multi-messenger targets for both electromagnetic (in radio domain) and gravitational wave astronomy. To outline the contours of a future physical model connecting SMBHB evolution with detectable GW manifestations, we present a simplified (‘toy’) model and, as an example, apply this toy model to the described in this work astrometrically oscillating source J02041514. We also provide justification for aiming future space-borne VLBI missions toward direct imaging of SMBHBs as a synergistic contribution into future multi-messenger studies involving prospective GW facilities.
@@
Formation of supermassive black hole binaries (SMBHBs) is one of the most challenging problems of theoretical astrophysics and observational electromagnetic and gravitational wave astronomy. Their formation and following evolution could be related to the merger of galaxies (Begelman et al., 1980), which might contain supermassive black holes (SMBHs) of any origin in their central regions. SMBHBs might be a result of merges of ‘ordinary’ black holes formed in the process of stellar or galactic evolution. The latter scenario has been analysed for one of the first discovered black hole binaries in the core of the object 3C 66 (De Paolis et al., 2004). Alternatively, black hole binary systems might be formed from primordial black holes (PBHs) born in the very early Universe (e.g., Polnarev et al., 2012; Marcoccia et al., 2024; Andrés-Carcasona et al., 2024; Escrivà et al., 2024, and references therein).
@@
However, regardless of the black hole binary genesis, they evolve and become ‘violent’ at the last stages of their orbital evolution (sometimes referred to as recoiling or inspiralling) before coalescing and emitting more and more appreciable gravitational waves (GW). Recently, predictions on the SMBHBs population were re-assessed by D’Orazio & Charisi (2023) and de la Parra et al. (2024). On the basis of detected sinusoidal flux density variations in two blazars, the latter work predicts that about 1% of blazars should contain SMBHB candidates.
@@
A pivotal direct detection of GW was made by the LIGO and Virgo collaborations in 2015 (Abbott et al., 2016). Although this discovery dealt with an act of coalescence of stellar mass black holes, it stimulated an outburst of interest in evolution of SMBHBs as generators of GW. The next powerful stimulus of interest in SMBHBs came recently with the announcement of detection of GW background by multiple Pulsar Timing Array (PTA) monitoring programs, in particular, by the North American Nanohertz Observatory for Gravitational Waves, NANOgrav (Agazie et al., 2023, and references therein). Very recent data by the James Webb Space Telescope (JWST) in comparison with the NANOGrav GW background measurements apparently favour scenarios in which the SMBHB are either ‘stalled’ in their evolution or undergo accelerated inspiralling in the presence of gas in host galaxies (Padmanabhan & Loeb, 2024). New quantitative evaluations of SMBHBs’ contribution in the GW background reported in the recent data releases by PTAs (Goncharov et al., 2024; Sato-Polito et al., 2025, and references therein). It should be noted however, that other origins of PTA GW stochastic background are possible, alternative to SMBHB coalescence (e.g., Braginsky et al., 1990; Winkler & Freese, 2025; Ferreira et al., 2024, and references therein).
@@
According to Begelman et al. (1980) (see also Colpi, 2014; De Rosa et al., 2019, and references therein), the formation and evolution of SMBHB encompasses three phases: the pairing phase, the hardening phase, and the final coalescence phase. The inhabitants of the SMBHB ‘zoo’, albeit small so far in numbers, are classified according to the primary evidence of the binary morphology, such as transiting, apparent, telescopic and astrometric binaries (see Ayzenberg et al., 2025, section 6.2 and Table 3). While the number of SMBHB candidates and corresponding publications grow very rapidly over the past several years, SMBHBs as a class of extragalactic objects remain rather elusive. The enigmatic apparent deficit of SMBHBs has become especially topical in view of the recent NANOGrav results (Sato-Polito et al., 2024). Some optimism about finding more binary active galactic nuclei (AGNs) comes from the recent ‘varstrometric’ (i.e., variability-induced astrometric jitter in the photocenter of the unresolved system, Hwang et al., 2020) analysis of a Gaia-based sample of quasars using Very Long Baseline Array (VLBA) observations (Chen et al., 2023). This study provides an indication that two out of 23 observed sources, J10442959 and J11103653, are possible binary quasars with the separation at the parsec scale. According to another rather optimistic prognosis, prospective mm/sub-mm VLBI studies (frequencies 86–690 GHz) will detect about 20 SMBHBs with masses at (Zhao et al., 2024). A forward look on LISA (Laser Interferometer Space Antenna) contribution into studies of SMBHBs as multi-messenger sources in the context of their evolution through the hardening phase is given recently by Spadaro et al. (2025).
@@
Arguably, one of the most ‘trusted’ SMBHB candidates is the BL Lacertae object OJ 287 (Britzen et al., 2018, and references therein). Its potential SMBHB nature has been corroborated by recent polarimetric Space Very Long Baseline Interferometry (SVLBI) observations by the RadioAstron mission at 22 GHz and nearly simultaneous Earth-based VLBI observations at 15, 43, and 86 GHz (Gómez et al., 2022). In this source, indications on the binary nucleus come from observations of different types, including an evidence of jet precession in VLBI images (Britzen et al., 2018) and optical flares (Dey et al., 2018). The estimates of the total mass of the binary system obtained with the VLBI and optical variability techniques show a notable discrepancy. The precession time scale analysed by Britzen et al. (2018) favours the total mass of the OJ 287 system of . The optical flare analysis leads to the total mass of (Dey et al., 2018, and references therein).
@@
A somewhat similar case of a suspected binary system containing an SMBH is reported in the nucleus of a galaxy associated with the flare ASASSN-20qc on the basis of its ‘after-flare’ quasi-periodic activity detected by multiple high-energy instruments (Pasham et al., 2024). In this case, the authors demonstrate indications of a binary system containing an SMBH and an intermediate-mass black hole (IMBH).
@@
Another recently reported suspected case of a sub-parsec-separated SMBHB is based on gravitational self-lensing manifestations detected in the Seyfert galaxy NGC 1566 (Kollatschny & Chelouche, 2024). The more massive component in this binary system responsible for the lensing effect has an estimated mass of , with a secondary object 10 times lighter.
@@
As another example, repeating major outbursts of the quasar J1048+7143 detected both in radio and -ray bands could be successfully modelled with the spin–orbit precession of the more massive jetted black hole with mass , tightly constraining the ratio of smaller black hole mass to the mass of larger BH: (Kun et al., 2022, 2024).
@@
To date, two SMBH components have been detected directly only in a few dual or binary systems. (Following the terminology of De Rosa et al. (2019), we define dual AGNs as not yet gravitationally bound accreting SMBH systems residing in the same host galaxy, and binary AGNs at sub-pc projected separation.) One of such systems is the radio galaxy 0402379 (Rodriguez et al., 2006; Bansal et al., 2017). In this source, monitored with multi-frequency VLBI imaging observations over 12 years, a pair of orbiting black holes with the total mass , identified in the milliarcsecond-scale morphology, demonstrates an orbital motion with the period yr and a projected separation between the components of 7.3 pc. In recent work, Surti et al. (2024) give a somewhat higher value of the nucleus’ mass in this object, (one of the heaviest black hole systems known), and argue that the pair of orbiting black holes in this source might be already gravitationally bound, thus constituting an SMBHB.
@@
A significantly larger apparent offset between two nuclei, pc, has been reported in the recently identified and dynamically confirmed case of the SMBHB system NGC 7277 (Voggel et al., 2022). The authors note that the smaller SMBH in this system is in an advanced state of inspiral, and the mass ratio of the two SMBHs in this system is . A comparable separation between SMBHB components of pc is also reported in the object UGC 4211 on the basis of Atacama Large Millimeter/submillimeter Array (ALMA) observations (Koss et al., 2023). A smaller separation of pc has been recently reported by Trindade Falcão et al. (2024) in the resolved dual nuclei in the infrared galaxy MCG-03-34-64 at on the basis of imaging and spectroscopic observations in radio (Very Large Array, VLA), optical (Hubble Space Telescope) and X-ray (Chandra, Suzaku, and XMM-Newton) bands. A new focused search for sub-kpc dual AGN based on HST and VLA imaging with several likely new canddates is presented by Gross et al. (2024).
@@
Several suspected dual AGN systems reported recently at kpc scales. Xu et al. (2024) investigated four such dual systems selected on the basis of SDSS and VLA observations. They emphasise that two systems with projected separations between dual AGNs between 7 and 8 kpc, contain components dubbed J00510020B and J23000005A with clear jet activity on pc-scales detected by VLBA. A new case of dual AGN with a separation of 1 kpc and mass ratio of black holes 7:1 is reported by Cao et al. (2025). This dual system is located within an edge-on disk galaxy, SDSS J14454926, and consists of two BHs of masses and . A case of two merging QSOs, HSC J121503.42014858.7 and HSC J121503.55014859 at one of the highest known to date redshift for such systems, , with the projected separation between components 12 kpc has been serendipitously discovered recently by the Subaru telescope (Matsuoka et al., 2024). It should be noted that the population of apparent double/triple quasars (AGNs) with separations of the order kpc is much more populous (see, e.g. Chen et al., 2022) than ‘likely’ SMBHBs separated by parsecs and closer. However, a somewhat opposite trend is reported recently by Saeedzadeh et al. (2024). They claim that the number of dual AGNs with separations 0.5–4 kpc is twice the number of duals with separations 4–30 kpc. This study also indicates an increase in the number of dual AGNs (as well as singles) from lower to higher redshift.
@@
To reveal the geometric and kinematic properties of SMBHBs, high angular resolution observational data are required. In this sense, VLBI comes as a natural technique for search and studies of potential SMBHB systems (e.g., An et al., 2018), provided that the AGNs are radio-emitting. This technique offers two measuring products, source images and their astrometric parameters. One of the convincing examples of SMBHB discovery using VLBI observations has been reported for the radio galaxy 3C 66B (Sudou et al., 2003). In that work, in-beam phase-referencing relative astrometry VLBI observations of the unresolved radio core in 3C 66B with respect to the background source 3C 66A resulted in the detection of a well defined elliptical motion of the former with a period of yr. A straightforward interpretation of such the kinematic pattern is a Keplerian motion in a binary system. Several more cases of astrometry-based suspected SMBHBs are mentioned in Ayzenberg et al. (2025, section 6.2.2 and references therein).
@@
Recently we have found that several radio sources observed in VLBI astrometry monitoring programs demonstrate a remarkable ‘oscillating’ behaviour of its celestial coordinates. Among others, these include J21026015 (2101600) (Titov et al., 2023; Gurvits et al., 2023) and J02041514 (0202149) presented in this work. We underline that these findings are based on absolute VLBI astrometric measurements and in this sense differ from the relative VLBI astrometric study of 3C 66B by Sudou et al. (2003) mentioned above. We consider the detected oscillations as a manifestation of a Keplerian motion of the SMBHB’s components as one of possible causes of the effect. Using this assumption, we further analyse their behaviour and evolution as potential contributors to the GW background and precursors of GW bursts. We note however that this is not a sole possible explanation of the apparent oscillations. The other plausible cause of the effect might be variations of the position of centroid of radio brightness distribution due to consecutive emission of jet components and their motion. We note that this explanation requires a separate study as a follow up on the work by Charlot (1990) and in general is of high importance for improving the precision of the International Celestial Reference Frame (ICRF, Charlot et al., 2020).
@@
In Section 2, we present the case of J02041415 as astrometrically oscillating source. We briefly describe its properties known from the literature (subsection 2.1) and analyse the VLBI astrometric measurements (subsection 2.2). The latter subsection contains description of statistical evaluation of the obtained parameters of the oscillations. Section 3 is dedicated to qualitative evaluation of the astrometric pattern of J02041415. In Section 4, we assume a ‘toy’ model of the sources in which the detected astrometric oscillations are considered as manifestations of SMBHBs. We further analyse orbital evolution of these SMBHBs and evaluate emitted by them gravitational waves, including potential contribution of these or similar objects in the GW background detected by pulsar timing arrays, as well as prospects of detection of GW bursts produced in final coalescence by SMBHBs. Section 5 provides a brief outlook of future search for SMBHBs as objects of multi-messenger studies. Section 6 lists several conclusions from the current study. In Appendix A, we present a new estimate of the redshift of the source J21026015, astrometric oscillations of which were reported by Titov et al. (2023).

### v2

# Milliarcsecond astrometric oscillations in active galactic nuclei as a precursor of multi-messenger gravitational wave events

@@ Abstract
Aims. Previously we have identified the quasar J21026015 as a potential SMBHB system on the basis of absolute astrometric very long baseline interferometry (VLBI) monitoring. Here we present another case, the source J02041514, exhibiting a similar oscillating astrometric pattern. Our aim is to analyse the evolution of SMBHBs as generators of GW and provide a physical ‘multi-messenger’ link between astrometric manifestation in the radio domain and GW emission.
Methods. We analysed the available archive VLBI astrometry data that resulted in the detection of astrometric oscillations in the source J02041514. We assume these oscillations to be manifestations of orbital motion in a binary system. We estimated the parameters of the suspected SMBHB in this source and applied basic theoretical models to project its evolution towards coalescence. We also developed a simplified ’toy’ model of SMBHBs consistent with the discovered astrometric oscillations and made quantitative predictions of GW emission of such sources using the case of J02041514 as an example potentially applicable to other SMBHBs.
Results. We provide observational evidence of astrometric oscillations in the source J02041514. As an ad hoc result, we also provide a re-assessed estimate of the redshift of J21026015, . A toy model of the object containing a SMBHB with parameters consistent with the observed astrometric oscillations of the source J02041514 as an example enables us to consider GW emission as the cause of the system’s orbital evolution.
Conclusions. We conclude that astrometric VLBI monitoring has an appreciable potential for future detections of SMBHBs that could become multi-messenger targets for both electromagnetic (in radio domain) and GW astronomy. To outline the contours of a future physical model connecting SMBHB evolution with detectable GW manifestations, we present a toy model and, as an example, apply this toy model to the astrometrically oscillating source J02041514 described in this work. We also provide a justification for aiming future space-borne VLBI missions at direct imaging of SMBHBs as a synergistic contribution into future multi-messenger studies involving prospective GW facilities.
@@
The formation of supermassive black hole binaries (SMBHBs) is one of the most challenging problems of theoretical astrophysics and observational electromagnetic and gravitational wave (GW) astronomy. Their formation and subsequent evolution could be related to the merger of galaxies (Begelman et al. 1980), which might contain supermassive black holes (SMBHs) of any origin in their central regions. Mergers of ‘ordinary’ black holes formed through stellar or galactic evolution may lead to the formation of SMBHBs.The latter scenario has been analysed for one of the first discovered black hole binaries in the core of the object 3C 66 (De Paolis et al. 2004). Alternatively, black hole binary systems might be formed from primordial black holes (PBHs) born in the very early Universe (e.g. Polnarev et al. 2012; Marcoccia et al. 2024; Andrés-Carcasona et al. 2024; Escrivà et al. 2024, and references therein).
@@
However, regardless of the black hole binary genesis, they evolve and become ‘violent’ at the last stages of their orbital evolution (sometimes referred to as recoiling or inspiralling) before coalescing and emitting more and more appreciable GWs. Recently, predictions about the SMBHB population were re-assessed by D’Orazio et al. (2024) and de la Parra et al. (2024). On the basis of sinusoidal flux density variations detected in two blazars, the latter work predicts that about 1% of blazars should contain SMBHB candidates.
@@
A pivotal direct detection of GW was made by the LIGO and Virgo collaborations in 2015 (Abbott et al. 2016). Although this discovery dealt with an act of coalescence of stellar-mass black holes, it stimulated an outburst of interest in the evolution of SMBHBs as generators of GW. The next powerful stimulus of interest in SMBHBs came recently with the announcement of the detection of a GW background by multiple pulsar timing array (PTA) monitoring programs, in particular by the North American Nanohertz Observatory for Gravitational Waves, NANOgrav (Agazie et al. 2023, and references therein). Very recent data from the James Webb Space Telescope (JWST) in comparison with the NANOGrav GW background measurements apparently favour scenarios in which the SMBHBs are either ‘stalled’ in their evolution or undergo accelerated inspiralling in the presence of gas in host galaxies (Padmanabhan & Loeb 2024). New quantitative evaluations of SMBHBs’ contribution to the GW background are reported in the recent data releases by PTAs (Goncharov et al. 2024; Sato-Polito et al. 2025, and references therein). It should be noted, however, that other potential sources of the stochastic GW background observed by PTAs may exist besides SMBHB coalescence (e.g. Braginsky et al. 1990; Winkler & Freese 2025; Ferreira et al. 2024, and references therein).
@@
According to Begelman et al. (1980) (see also Colpi 2014; De Rosa et al. 2019, and references therein), the formation and evolution of SMBHB encompasses three phases: the pairing phase, the hardening phase, and the final coalescence phase. The inhabitants of the SMBHB ‘zoo’, albeit small so far in numbers, are classified according to the primary evidence of the binary morphology, such as transiting, apparent, telescopic, and astrometric binaries (see Ayzenberg et al. 2025, section 6.2 and Table 3). While the number of SMBHB candidates and corresponding publications have grown very rapidly over the past several years, SMBHBs as a class of extragalactic objects remain rather elusive. The enigmatic apparent deficit of SMBHBs has become especially topical in view of the recent NANOGrav results (Sato-Polito et al. 2024). Some optimism about finding more binary active galactic nuclei (AGNs) comes from the recent ‘varstrometric’ (i.e. variability-induced astrometric jitter in the photocenter of the unresolved system, Hwang et al. 2020) analysis of a Gaia-based sample of quasars using Very Long Baseline Array (VLBA) observations (Chen et al. 2023). This study provides an indication that two out of 23 observed sources, J10442959 and J11103653, are possible binary quasars with the separation at the parsec scale. According to another rather optimistic prognosis, prospective millimetre/sub-millimetre VLBI studies (frequencies 86–690 GHz) will detect about 20 SMBHBs with masses of at (Zhao et al. 2024). A forward look at the contribution of LISA (Laser Interferometer Space Antenna) to studies of SMBHBs as multi-messenger sources in the context of their evolution through the hardening phase has recently been given by Spadaro et al. (2025).
@@
Arguably, one of the most ‘trusted’ SMBHB candidates is the BL Lacertae object OJ 287 (Britzen et al. 2018, and references therein). Its potential SMBHB nature has been corroborated by recent polarimetric space very long baseline interferometry (SVLBI) observations by the RadioAstron mission at 22 GHz and nearly simultaneous Earth-based VLBI observations at 15, 43, and 86 GHz (Gómez et al. 2022). In this source, indications of the binary nucleus come from observations of different types, including evidence of jet precession in VLBI images (Britzen et al. 2018) and optical flares (Dey et al. 2018). The estimates of the total mass of the binary system obtained with the VLBI and optical variability techniques show a notable discrepancy. The precession time scale analysed by Britzen et al. (2018) favours the total mass of the OJ 287 system of . The optical flare analysis leads to the total mass of (Dey et al. 2018, and references therein).
@@
A somewhat similar case of a suspected binary system containing an SMBH is reported in the nucleus of a galaxy associated with the flare ASASSN-20qc on the basis of its ‘after-flare’ quasi-periodic activity detected by multiple high-energy instruments (Pasham et al. 2024). In this case, the authors demonstrate indications of a binary system containing an SMBH and an intermediate-mass black hole (IMBH).
@@
Another recently reported suspected case of a sub-parsec-separated SMBHB is based on gravitational self-lensing manifestations detected in the Seyfert galaxy NGC 1566 (Kollatschny & Chelouche 2024). The more massive component in this binary system responsible for the lensing effect has an estimated mass of , with a secondary object ten times lighter.
@@
As another example, repeating major outbursts of the quasar J1048+7143 detected both in radio and -ray bands could be successfully modelled with the spin–orbit precession of the more massive jetted black hole with mass , tightly constraining the ratio of the smaller black hole mass, , to the mass of the larger BH: (Kun et al. 2022, 2024).
@@
To date, two SMBH components have been directly detected in only a few dual or binary systems. (Following the terminology of De Rosa et al. (2019), we define dual AGNs as not yet gravitationally bound accreting SMBH systems residing in the same host galaxy, and binary AGNs at sub-parsec projected separation.) One such system is the radio galaxy 0402379 (Rodriguez et al. 2006; Bansal et al. 2017). In this source, monitored with multi-frequency VLBI imaging observations over 12 years, a pair of orbiting black holes with the total mass , identified in the milliarcsecond-scale morphology, demonstrates an orbital motion with the period yr and a projected separation between the components of 7.3 pc. In a recent work, Surti et al. (2024) give a somewhat higher value of the nucleus’ mass in this object, (one of the heaviest black hole systems known), and argue that the pair of orbiting black holes in this source might already be gravitationally bound, and thus constitute an SMBHB.
@@
A significantly larger apparent offset between two nuclei, pc, has been reported in the recently identified and dynamically confirmed case of the SMBHB system NGC 7277 (Voggel et al. 2022). The authors note that the smaller SMBH in this system is in an advanced state of inspiral, and the mass ratio of the two SMBHs in this system is . A comparable separation between SMBHB components of pc is also reported in the object UGC 4211 on the basis of Atacama Large Millimeter/submillimeter Array (ALMA) observations (Koss et al. 2023). A smaller separation of pc has recently been reported by Trindade Falcão et al. (2024) in the resolved dual nuclei in the infrared galaxy MCG-03-34-64 at on the basis of imaging and spectroscopic observations in radio (Very Large Array, VLA), optical (Hubble Space Telescope), and X-ray (Chandra, Suzaku, and XMM-Newton) bands. A new focused search for sub-kiloparsec dual AGNs based on HST and VLA imaging with several likely new candidates is presented by Gross et al. (2024).
@@
Several suspected dual AGN systems have been reported recently at kiloparsec scales. Xu et al. (2024) investigated four such dual systems selected on the basis of SDSS and VLA observations. They emphasise that two systems with projected separations between dual AGNs of between 7 and 8 kpc contain components dubbed J00510020B and J23000005A with clear jet activity on parsec scales detected by VLBA. A new case of dual AGNs with a separation of 1 kpc and black hole mass ratio of 7:1 is reported by Cao et al. (2025). This dual system is located within an edge-on disc galaxy, SDSS J14454926, and consists of two BHs of masses and . A case of two merging QSOs, HSC J121503.42014858.7 and HSC J121503.55014859 at one of the highest redshifts known to date for such systems, , with a projected separation between the components of 12 kpc, was recently serendipitously discovered by the Subaru telescope (Matsuoka et al. 2024). It should be noted that the population of apparent double or even triple quasars (AGNs) with separations of the order kpc is much more populous (see, e.g. Chen et al. 2022) than ‘likely’ SMBHBs separated by parsecs and closer. However, a somewhat opposite trend has recently been reported by Saeedzadeh et al. (2024). They claim that the number of dual AGNs with separations of 0.5–4 kpc is twice the number of duals with separations of 4–30 kpc. This study also indicates an increase in the number of dual AGNs (as well as singles) from lower to higher redshifts.
@@
To reveal the geometric and kinematic properties of SMBHBs, high-angular-resolution observational data are required. In this sense, VLBI comes as a natural technique in searching for and studying potential SMBHB systems (e.g. An et al. 2018), provided that the AGNs are radio-emitting. This technique offers two measuring products: source images and their astrometric parameters. One convincing example of SMBHB discovery using VLBI observations was reported for the radio galaxy 3C 66B (Sudou et al. 2003). In that work, in-beam phase-referencing relative astrometry VLBI observations of the unresolved radio core in 3C 66B with respect to the background source 3C 66A resulted in the detection of the well-defined elliptical motion of the former with a period of yr. A straightforward interpretation of such a kinematic pattern is Keplerian motion in a binary system. Several more cases of astrometry-based suspected SMBHBs are mentioned in Ayzenberg et al. (2025, section 6.2.2 and references therein).
@@
Recently we have found that several radio sources observed in VLBI astrometry monitoring programs demonstrate a remarkable ‘oscillating’ behaviour in their celestial co-ordinates. Among others, these include J21026015 (2101600) (Titov et al. 2023; Gurvits et al. 2023) and J02041514 (0202149) presented in this work. We underline that these findings are based on absolute VLBI astrometric measurements and in this sense differ from the relative VLBI astrometric study of 3C 66B by Sudou et al. (2003) mentioned above. We consider a Keplerian motion of the SMBHB’s components as one of possible causes of the detected astrometric oscillations. Using this assumption, we further analyse their behaviour and evolution as potential contributors to the GW background and precursors of GW bursts. We note, however, that this is not a sole possible explanation of the apparent oscillations. The other plausible cause of the effect might be variations in the position of the centroid of radio brightness distribution due to consecutive emission of jet components and their motion. We note that this explanation requires a separate study as a follow-up on the work by Charlot (1990) and in general is of high importance for improving the precision of the International Celestial Reference Frame (ICRF, Charlot et al. 2020).
@@
In Section 2, we present the case of J02041415 as an astrometrically oscillating source. We briefly describe its properties known from the literature (subsection 2.1) and analyse the VLBI astrometric measurements (subsection 2.2). The latter subsection contains a description of the statistical evaluation of the obtained parameters of the oscillations. Section 3 is dedicated to qualitative evaluation of the astrometric pattern of J02041415. In Section 4, we assume a toy model of the sources in which the detected astrometric oscillations are considered as manifestations of SMBHBs. We further analyse the orbital evolution of these SMBHBs and evaluate the GWs emitted by them, including the potential contribution of these or similar objects in the GW background detected by PTAs, as well as prospects of detecting GW bursts produced in final coalescence by SMBHBs. Section 5 provides a brief outlook of a future search for SMBHBs as objects of multi-messenger studies. Section 6 lists several conclusions from the current study. In Appendix A, we present a new estimate of the redshift of the source J21026015, astrometric oscillations of which were reported by Titov et al. (2023).